Antonio rides his bike 7.3 miles to school. He wants to change his route so that he rides 4 miles less than twice the distance he is currently riding. Choose the equation that would be used to find the number of miles, m, of Antonio's new route.
A. 4-2m=7.3
B. 7.3-m=4
C. 2m-7.3=4
D. 2(7.3)-4=m

Answer by Alan3354(69443)

Antonio rides his bike 7.3 miles to school. He wants to change his route so that he rides 4 miles less than twice the distance he is currently riding. Choose the equation that would be used to find the number of miles, m, of Antonio's new route.
A. 4-2m=7.3
B. 7.3-m=4
C. 2m-7.3=4
D. 2(7.3)-4=m
==================================
He's going 7.3 miles now, so it'll be 2*7.3, right?